40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 935     Entry time: Mon Sep 8 10:57:49 2008
Author: steve 
Type: Update 
Category: IOO 
Subject: the psl and mc are back to normal 
The alarm handler is silent this morning.
This is almost unbelievably pleasant after two mount of harassment.
The MC did not lose lock for three days.

Atm1: the new fss layout
Atm2: PMC with lead brick
Atm3: 3 days plot
Attachment 1: fss.png  1.204 MB  | Hide | Hide all
fss.png
Attachment 2: pmcbrick.png  368 kB  | Hide | Hide all
pmcbrick.png
Attachment 3: brick.jpg  95 kB  Uploaded Mon Sep 8 12:20:37 2008  | Hide | Hide all
brick.jpg
ELOG V3.1.3-